fred is double as old as his younger brother philip but only half as old as his older sister anna. their mother will soon be 40 years old, and then double as old as anna. how old are the kids
Answer by T-Maree(17)

Anna = 20
Fred = 10
Philip = 5
40/2 = 20 =Anna's age
20/2 = 10 =Fred's age
10/2 = 5 =Philip's age